Flame Holdings is a company with a called up and paid up capital of 100,000 ordinary

shares of £1 each and 20,000 10% redeemable preference shares of £1 each.

The gross profit was £200,000 and trading expenses were £50,000. Flame Holdings

paid the required preference share dividend and an ordinary dividend of £ 0.42 per

share. The tax charge for the year was estimated at £40,000.

Required

Calculate basic EPS for the year


1
Expert's answer
2022-03-07T18:07:26-0500

"Basic EPS = \frac{(Net income - preferred dividends)}{ weighted average of common shares outstanding during the period}"


Net income=200 000-50 000-40 000=110 000


"preferred dividends=20 000\times0.42=8400"


"Basic EPS = \frac{(110 000- 8400)}{100000}=1.016"